OpenStudy (anonymous):

solve the system by the elimination method. 5x+3y=-11 7x-2y=17

OpenStudy (anonymous):

So, am I to calculate like this? S.A. of sphere= 4*3.14*r^2=4*3.14*17.5*17.5=3846.5 approx to 3850. which is option(b)?

OpenStudy (anonymous):

I am having so much trouble with these. I am not sure.

OpenStudy (anonymous):

multiply first equation by 2 and second by 3 Then add the second equation to first and x will be eliminated, n u get value of y then u can put this value of y in first equation to get value of x.....

OpenStudy (anonymous):

Elimination works by combining both equations and eliminating one of the variables in the process. First, since you can multiply both sides of the equal sign and maintain the validity of your equation, multiply the top equation by 2 and the bottom one by 3 to get: 10x+6y=-22 21x-6y=51 Now just add both equations and the y variable gets eliminated and you get one equation with one variable. Then just solve for x. Once you have the value for x, just substitute into either equation and you get one equation with one variable, then just solve for y and you're done.

OpenStudy (anonymous):

try it and tell me what u get.... if u want u can post each step and i will confirm it... this way u will learn how to do more sums....

OpenStudy (anonymous):

sorry I made a mistake ... om doing the steps i suggested , you will eliminate y and get value of x...

OpenStudy (anonymous):

I understand the first part, but still fuzzy on how to do the second part, and i have like 8 problems like this.. Help Please!!!

jhonyy9 (jhonyy9):

5x +3y = -11 / 2 --- 10x +6y = -22 7x -2y = 17 / 3 --- 21x -6y = 51 -------------- 31x = 29 --- x=29/31

OpenStudy (anonymous):

no I tryed that it is not right is that its simplist form

jhonyy9 (jhonyy9):

ok but this is the elimination method

OpenStudy (anonymous):

what are the integers

jhonyy9 (jhonyy9):

so if x =29/31 --- y = (-22 -290/31)/6 = (-682 -290)/186 = -972/186 = = - 486/93 x=29/31 y= -972/186

jhonyy9 (jhonyy9):

sorry y= - 486/93